OpenStudy (anonymous):

calc 3 help!!! use a triple integral to find the volume of the solid enclosed by the sphere x^2+y^2+z^2=2a^2 and the paraboloid az = x^2+y^2 (a>0)

OpenStudy (anonymous):

first find the point of intersection of surfaces since x^2+y^2=az put in sphere equation az+z^2=2a^2 z^2+az-2a^2=0 solve above equation so z=a (from given situation the other root is neglected) now using z=a in any of the above given equations of surfaces x^2+y^2=a^2 using cylindrical coordinates since x^2+y^2=r^2 so

OpenStudy (anonymous):

i hope you can do this integral if not let me know :)

OpenStudy (anonymous):

@mickifree12 can you do this ?

OpenStudy (anonymous):

not really, looking at this problem and a spherical coordinates one

OpenStudy (anonymous):

would I treat a as a constant o.0

OpenStudy (anonymous):

yes it is just a radius

OpenStudy (anonymous):

I'm stuck =[

OpenStudy (anonymous):

where ? in the integral? just integrate first with respect to z treat r as constant then integrate with respect to r (now treat it as variable) and then integrate with respect to thetha.

OpenStudy (anonymous):

wait, I see what you did now, I was making your "plug az into az+z^2=2a^2" more complicated than it really was

OpenStudy (anonymous):

I can't solve the z^2+az-2a^2=0 to the point where you got z=a

OpenStudy (anonymous):

it is just quadratic equation let me solve it

OpenStudy (anonymous):

oohhh that's why I didn't see it, I rarely use the quadratic equation to the point where I don't even know it my memory xP

OpenStudy (anonymous):

ok here we have

OpenStudy (anonymous):

got it now?

OpenStudy (anonymous):

yes, I wasn't sure if I solve for Z or A, but... where are you getting these pictures from o.0?

OpenStudy (anonymous):

i am typing in software known as Mathtype used for typing Maths .then take picture and upload here :)

OpenStudy (anonymous):

aaawwww, I thought you somehow had a pdf of a book and it was done out in the book =[[

OpenStudy (anonymous):

i use it because it is easy to use and Maths is readable like textbook format!

OpenStudy (anonymous):

why are you neglecting that z=-2a?

OpenStudy (anonymous):

look at the picture below

OpenStudy (anonymous):

since it is bounded by sphere above and paraboloid below taking z=-2a we are going away from the reigion of our interest.

OpenStudy (anonymous):

ok, so from x^2+y^2=r^2

OpenStudy (anonymous):

would I make the x and y cos and sin?

OpenStudy (anonymous):

do not need to do this will get complex .

OpenStudy (anonymous):

we just needs bounds for r and thetha in polar integrals .

OpenStudy (anonymous):

I thought when you convert to r and theta, you convert your x and y into cos and sin that way when you have something like x^2+y^2, when you convert them, it becomes simply 1

OpenStudy (anonymous):

I don't exactly know spherical/cylindrical coordinates that well since my teacher just went over them

OpenStudy (anonymous):

read these articles in your textbook.these are important.

OpenStudy (anonymous):

did u solve the integral?

OpenStudy (anonymous):

No, still trying to figure out this cylindrical coordinates right now

OpenStudy (anonymous):

How did you know that d theta was from 0-2pi?

OpenStudy (anonymous):

because x^2+y^2=a^2 is circle of radius r and for circle thetha goes from 0 to 2pi.

OpenStudy (anonymous):

aaahh, doing the integral right now, just messed up and did the first integration with respect to r instead of z, reworking it now

OpenStudy (anonymous):

for the final answer, would a still be present o.0?

OpenStudy (anonymous):

yes it will be

OpenStudy (anonymous):

I'm getting a super ugly integration right when I do my integration with respect to r

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

ok, just making sure I'm not completely off

OpenStudy (anonymous):

is it....

OpenStudy (anonymous):

[2a^2/3 - 1/3(2a^2-81)^(3/2) - (729/4)]2pi

OpenStudy (anonymous):

no you made some mistake check it !:)

OpenStudy (anonymous):

ggaarrrrrrr

OpenStudy (anonymous):

after the first integration...

OpenStudy (anonymous):

rsqrt(2a^2-r^2) - r^3/4?

OpenStudy (anonymous):

after first integration u get rsqrt(2a^2-r^2) - r^3/a

OpenStudy (anonymous):

yeah oops, accidentally wrote 4, instead of A

OpenStudy (anonymous):

and then.... 2nd integration, before plugging in 0 and 9

OpenStudy (anonymous):

-1/3 (2a^2-r^2)^(3/2) - r^4/4a

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

after plugging in

OpenStudy (anonymous):

here's where it gets ugly

OpenStudy (anonymous):

-1/3 (2a^2-81)^(3/2) - 6561/36) - (-1/3 (2a^2))

OpenStudy (anonymous):

after pluging in you should have 1/12(8sqrt(2)-7)a^3

OpenStudy (anonymous):

ok let me do it

OpenStudy (anonymous):

I'm currently redoing it, from that point

OpenStudy (anonymous):

I'm still getting the same answer, unless you somehow did some factoring o.0

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

@mickifree12

OpenStudy (anonymous):

for dr is it 0-a o.0

OpenStudy (anonymous):

this whole time I thought it said 0 to 9 -.-

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh jeez ahahahah

OpenStudy (anonymous):

i hope you can do it now :) best of luck for rest of ur studies.!